OpenStudy (anonymous):

What is the solution to this system of equations? 6x-2y=10 4x=-2y+5

OpenStudy (radar):

write second equation like this 4x+2y=5 then add the first equation 6x-2y=10 ____________ 10x=15

OpenStudy (radar):

solve for x and substitute x in one of the equation and solve for y

OpenStudy (radar):

Have you reached a solution yet?

OpenStudy (gw2011):

Rearrange the second equation and you will get the following equations: 6x-2y=10 4x+2y=5 Solve to get: 10x=15 x=15/10=3/2 Substitute the value of x into one of the equations to get y: 4(3/2)+2y=5 6+2y=5 2y=-1 y=-1/2 Therefore, x=3/2 and y=-1/2 You can substitute these values into one or both of the equations to prove the correctness of these values.

OpenStudy (anonymous):

yes, radar, i did. (: i just hadn't posted it, sorry. /:

OpenStudy (anonymous):

and thanks, gw2011! (:

OpenStudy (radar):

O.K, it is a good idean as gw2011 suggested to check by plugging your solutions in and checking the results.

OpenStudy (radar):

Good luck with your studying effort.

OpenStudy (anonymous):

thanks. (: but most of the problems on this homework are material i haven't seen since middle school... so i'm having to consult my book & google constantly.

OpenStudy (radar):

Understand, I too am trying to learn. Good bye